Carlson Auto Dealers Inc. sells a handmade automobile as its only product. Each automobile is identical; however, they can be distinguished by their unique ID number. At the beginning of 2018, Carlson had three cars in inventory, as follows:

Car ID Cost
203 $ 69,000
207 69,000
210 72,000


During 2018, each of the three autos sold for $99,000. Additional purchases (listed in chronological order) and sales for the year were as follows:

Car ID Cost Selling Price
211 $ 69,000 $ 99,000
212 69,000 102,000
213 70,500 not sold
214 72,000 105,000
215 75,000 109,500
216 73,500 not sold
217 78,000 114,000
218 75,300 115,500
219 81,000 not sold


Required:
1. Calculate 2018 ending inventory and cost of goods sold assuming the company uses the specific identification inventory method.
2. Calculate ending inventory and cost of goods sold assuming FIFO and a periodic inventory system.
3. Calculate ending inventory and cost of goods sold assuming LIFO and a periodic inventory system.
4. Calculate ending inventory and cost of goods sold assuming the average cost method and a periodic inventory system.

Specific Identification FIFO Periodic LIFO Periodic Average Cost Method Periodic
Ending inventory
Cost of goods sold
Answer #1

Sol:

Car ID Cost $
203 69,000
207 69,000
210 72,000
211 69,000
212 69,000
213 70,500
214 72,000
215 75,000
216 73,500
217 78,000
218 75,300
219 81,000
Total cost of goods available for sale ( total 12 cars ) 1026600

1. Specific identification inventory method :

Car ID 213, 216 and 219 is unsold, it indicates that these cars are in the ending inventory

Ending inventory = $70,500 + $73,500 + $81,000 = $225,000

Cost of goods sold = Total cost of goods available for sale - Ending inventory cost = $1026600 - $225,000 = $801,600

2. FIFO and a periodic inventory system :

In periodic inventory system, the transaction relating to the inventory are not updated immediately, but the same are updated at the end of a particular period.

In FIFO method those inventories are sold out first which are purchased first. It indicates that the ending inventory will be from the recent purchases.

Total 3 cars are unsold that will be in the ending inventory

Ending inventory = $81,000 + $75,300 + $78,000 = $234,300

Cost of goods sold = Cost of goods available for sale - ending inventory cost = $1026600 - $234,300 = $792,300

3. LIFO and a periodic inventory system :

In LIFO method those inventories are sold out first, which are purchased last i.e. purchased recently. It indicates that the ending inventory will be from the opening inventory and earlier purchases

Total 3 cars are unsold that will be in the ending inventory

Ending inventory = $69,000 + $69,000 + $72,000 = $210,000

Cost of goods sold = Cost of goods available for sale - ending inventory cost = $1,026,600 - $210,000 = $816,600

4. Average cost and a periodic inventory system :

Average cost per unit = cost of goods available for sale / units available for sale = $1,026,600 / 12 cars = $85,550per unit/car

Ending inventory = Average cost per unit * Ending inventory units = $85,550 * 3 cars = $256,650

Units sold = units available for sale - ending inventory units = 12 - 3 = 9

Cost of goods sold = Average cost per unit * Units sold = $85,550 * 9 cars = $769,950

Specific Identification FIFO Periodic LIFO periodic Average Cost Method Periodic
Ending inventory $225,000 $234,300 $210,000 $256,650
Cost of goods sold $801,600 $792,300 $816,600 $769,950